# Region-aware Grasp Framework with Normalized Grasp Space for 6-DoF Grasping in Cluttered Scene

### v1

@@ VI-A Cluttered Scene Clearance
To conduct a comprehensive comparison of grasp detection methods in real-world cluttered scenes, we constructed multiple cluttered scenes featuring 6 to 9 randomly placed objects within a $50cm\times 50cm$ workspace, including object stacking and occlusion, as shown in Fig. 9. We deployed multiple representative grasp detection methods [3, 10] on our real-world robot platform. Subsequently, akin to the metrics utilized in [3], we assessed the overall Success Rate (#$\#Success\_Attempts/\#All\_Attempts$) and Completion Rate (#$\#Cleared\_Scenes/\#All\_Scenes$) to facilitate overall comparison. More experiment settings can be found in the Supplementary Materials.
@@
In comparison with other methods, the results presented in Table VII demonstrate the generalization capability to real-world grasping of our framework. However, it is worth noting that HGGD struggles to generate grasps for objects with very small scales, and AnyGrasp exhibits imprecise rotation predictions, resulting in unstable object grasping and frequent drops, particularly evident when objects are stacked. Both methods are prone to failure when the captured point cloud becomes more noisy or only very partial object observations are available. Conversely, RNGNet exhibits greater robustness to noise originating from the depth sensor and demonstrates the ability to predict more stable grasps with precise rotations, thereby effectively reducing collisions and object drops.
VI-B Grasp Inference with Multi-scale Grippers
As clarified in Section. IV-B3, the Normalized Grasp Space (NGS) demonstrates scale invariance. In other words, our grasp detection model can be easily scaled to accommodate a different range from the fixed gripper width defined in the training dataset without any extra training. In Fig. 10, we illustrate this by enlarging the gripper scale from $w_{gripper}=0.05\ m$ to $w_{gripper}=0.2\ m$, showcasing grasps in a cluttered scene. These results affirm that our framework can seamlessly adapt to grippers of varying sizes. As is shown in Fig. 11, we extend the application of our framework to a much larger indoor scene than the typical table-top settings. The proposed RNGNet successfully generates large-scale grasps for large objects, validating the effectiveness of the Normalized Grasp Space.
To further investigate the effectiveness of our Normalized Grasping Space for object grasping with grippers of different scales, We build scenes with objects with significantly different scales in the Sapien [43] simulator and conduct grasp detection and robot grasping. Experiments prove that our Normalized Grasping Space can generalize well to grippers with different scales and has the potential to apply to industrial scenes, in which data and labels are hard to obtain. Detailed results and visualization can be found in our Supplementary Materials.
Figure 10: Visualizations of grasp predictions of different gripper scales in a real clutter scene. Different colors mean different gripper widths. From (a) to (e), as the gripper scale enlarges, it shows adaptive grasp predictions under different configurations.
Figure 11: Process for our method to detect grasp with an assigned large specific gripper width, $w_{gripper}=0.4\ m$. (a) Target RGB image and depth image of an ordinary indoor scene with a sofa placed. (b) Extracted patch RGB and normalized z image. (c) Detected grasps of the sofa.
Figure 12: Experiment settings and examples of our closed-loop grasping implemented in the real dynamic scenes.
VI-C Region-aware Closed-loop Grasping
With pretrained RNGNet, provided the corresponding patch, we are able to detect grasps in any region through the Normalized Grasp Space, regardless of the patch location or the scene structure. This implies a useful downstream application potential in specific scenarios when there is no need to generate a grasp all across the scene. Our region-aware framework only focuses on regions of interest, alleviating reliance on a specific grasp location model.
Therefore, we conduct trials to explore the application of our method in the downstream task of dynamic grasping in a more complicated scene, which can benefit from the efficiency and flexibility of our framework. As demostrated in the Algorithm. 1, we only need to integrate grasp detection with simple robot controlling and keep tracking the real-time grasp detection results to achieve closed-loop grasping in dynamic scenes.
@@ VI-C Region-aware Closed-loop Grasping
To evaluate the efficiency of our framework in more challenging scenes, we build two dynamic settings as illustrated in Fig. 12: (a) Human-to-robot handover; (b) Moving object on the conveyor. Experiment results in Table VIII confirm that our framework can accomplish human-to-robot handover and grasping the moving object on the conveyor at a satisfactory success rate by promptly generating a diverse range of high-quality grasps in a real-time frame rate and executing the closed-loop grasping (at around 10 FPS, due to extra time cost for robot controlling and image capturing). The proposed closed-loop grasping pipeline can effectively handle with more complex nontable-top grasping scenarios. Thus, our proposed method proves robust to much noisier environment, which underscores the potential of our method for addressing challenging dynamic scenarios. More demonstrations of the dynamic grasping experiments can be found in the Supplementary Materials.
However, we believe that the full potential for a robust patch-based grasp detection method remains to be fully explored and is worth deeper investigation.
@@
2:  $grasps=\text{list}()$
3:  ## Tracking phrase
4:  while $\text{dist}(grasps[0].\text{pose},robot.\text{pose})>min\_dis$ do
@@
6:     $centers=\text{uniform\_sample}(M_{workspace})$ sample patch centers within the mask
7:     $patches=\text{patch\_extract}(centers,image)$ $patches=\text{normalize}(patches)$ extract and normalize patch regions
8:     $grasps=\text{RNGNet}(patches)$ detect grasps for objects in the workspace
[…]
10:     $robot.\text{set\_speed}(grasps[0].\text{pose}-robot.\text{pose})$ robot moving towards the best grasp pose
11:  end while
12:  ## Grasping phrase
13:  $robot.\text{move\_pose}(grasps[0].\text{pose})$ $robot.\text{grasp}()$ move robot to the final grasp pose and grasp
TABLE VIII: Extensive Results for Region-aware Closed-loop Grasping
Method Scenario Success Rate
Closed-loop Human-to-Robot Handover 85% (17 / 20)
Moving Object on Conveyor 65% (13 / 20)
VII Limitations
@@
VIII Conclusion and Future Work
In this paper, we design a novel Region-aware Grasp Framework for 6-DoF grasp detection in cluttered scenes. Normalized Grasp Space (NGS), a shared unified space projected on the RGBD image plane for region-based grasp detection, is proposed to obtain consistent regional grasp representation. With the representation, we further build an efficient convolution-based Region Normalized Grasp Network (RNGNet) to extract regional features and predict regional grasps. Experiments on the dataset and real-world trials both prove the efficacy of our framework and indicate promising downstream applications. We plan to explore integrating scene-level information from the multi-view image during the process of closed-loop grasping, which is critical for precise manipulation in complex scenes or with specific instructions.

### v2

@@ 5 Realworld Experiments
Cluttered Scene Clearance: To conduct a comprehensive comparison of grasp detection methods in real-world cluttered scenes, we constructed multiple cluttered scenes consisting of 6 to 9 randomly placed objects within a $50cm\times 50cm$ workspace, including novel objects, object stacking and occlusion, as shown in Fig. 7. We deployed multiple representative grasp detection methods [15, 9] on our real-world robot platform. Subsequently, akin to the metrics utilized in [15], we assessed the overall Success Rate and Completion Rate to facilitate overall comparison.
@@
In comparison with other methods, the results presented in Table 5 demonstrate the generalization capability to real-world grasping of our framework. However, it is worth noting that HGGD struggles to generate grasps for objects with very small scales, and AnyGrasp exhibits imprecise rotation predictions, resulting in unstable object grasping and frequent drops, particularly evident when objects are stacked. Both methods are prone to failure when the captured point cloud becomes more noisy or only very partial object observations are available. In contrast, RNGNet, with its greater robustness to noise originating from the depth sensor, demonstrates the ability to predict more stable grasps with precise rotations. This effectively reduces collisions and object drops, showcasing the potential of this method.
Figure 8: Real-world closed-loop experiment settings.
Setting Success Rate
Handover 82.5%
(33 / 40)
Conveyor 67.5%
(27 / 40)
Table 6: Closed-loop results.
@@
Closed-loop Grasping: With RNGNet, we are able to detect grasps in any region regardless of the patch location or the scene structure. This implies a practical downstream application potential when there is no need to generate grasps across the scene. To evaluate the efficiency of our framework in more challenging scenes, we build two dynamic settings as illustrated in Fig. 8: (a) Human-to-robot handover; (b) Moving object on the conveyor. Experiment results in Table 6 confirm that our framework can accomplish dynamic grasping at a satisfactory success rate by detecting high-quality grasps in a real-time frame rate and executing the closed-loop controlling. The proposed closed-loop grasping pipeline can effectively handle more complex nontable-top grasping scenarios. Thus, our proposed method proves robust to a much noisier environment, which underscores the potential of our method for addressing challenging dynamic scenarios. The closed-loop grasping algorithm is demonstrated in Algorithm. 1 of the Appendices, and more demonstrations of the dynamic grasping experiments can be found in the Supplementary Videos.
6 Conclusion
We design a novel Region-aware Grasp Framework for 6-DoF grasp detection. Normalized Grasp Space (NGS), a unified space for region-based grasp detection, is proposed to obtain consistent region representation. With the representation, we further build an efficient convolution-based Region Normalized Grasp Network (RNGNet) to extract features and predict rotation heatmap. We plan to explore integrating scene-level information from the multi-view image during the process of closed-loop grasping, which is critical for precise manipulation in complex scenes or with specific instructions. More limitations are discussed in the Section C of the Appendices.
Appendix A Additional Results
Detailed Results on the GraspNet-1Billion Dataset: In addition to the overall grasp AP (Average Precision) metric, we also provide the evaluated grasp quality based on force-closure metric [27] with varying friction coefficients, denoting $AP_{\mu}$, representing the average precision at given friction coefficient $\mu$. As is shown in Table 7, our proposed RNGNet outperforms previous state-of-the-art methods in terms of all metrics under different friction conditions, especially on the difficult scenarios (similar/novel splits and $AP_{0.4}$ with lower friction coefficient).
Method Seen Similar Novel
$AP$ $AP_{0.8}$ $AP_{0.4}$ $AP$ $AP_{0.8}$ $AP_{0.4}$ $AP$ $AP_{0.8}$ $AP_{0.4}$
GPD [10] 22.87/24.38 28.53/30.16 12.84/13.46 21.33/23.18 27.83/28.64 9.64/11.32 8.24/9.58 8.89/10.14 2.67/3.16
PointnetGPD [11] 25.96/27.59 33.01/34.21 15.37/17.83 22.68/24.38 29.15/30.84 10.76/12.83 9.23/10.66 9.89/11.24 2.74/3.21
GraspNet-baseline [3] 27.56/29.88 33.43/36.19 16.95/19.31 26.11/27.84 34.18/33.19 14.23/16.62 10.55/11.51 11.25/12.92 3.98/3.56
TransGrasp [26] 39.81/35.97 47.54/41.69 36.42/31.86 29.32/29.71 34.80/35.67 25.19/24.19 13.83/11.41 17.11/14.42 7.67/5.84
HGGD [15] 58.35/56.85 66.54/64.60 55.96/52.94 47.93/43.93 56.91/52.73 41.86/36.88 22.10/18.59 27.37/22.98 14.31/11.91
Scale Balanced Grasp [8] 58.95/  - 68.18/  - 54.88/  - 52.97/  - 63.24/  - 46.99/  - 22.63/  - 28.53/  - 12.00/  -
GSNet [6] 65.70/61.19 76.25/71.46 61.08/56.04 53.75/47.39 65.04/56.78 45.97/40.43 23.98/19.01 29.93/23.73 14.05/10.60
AnyGrasp w/ CD [9] 66.12/  - 77.27/  - 61.02/  - 56.09/  - 68.29/  - 47.31/  - 24.81/  - 31.08/  - 13.82/  -
RNGNet 75.20/72.23 85.02/82.04 71.71/67.95 66.62/58.43 78.53/68.69 59.23/51.87 32.38/26.05 40.44/32.29 19.84/16.24
RNGNet w/ CD 76.28/72.89 86.58/83.10 72.26/68.11 68.26/59.42 80.86/70.16 60.01/52.13 32.84/26.12 41.07/32.45 19.68/15.92
“-”: Result Unavailable, CD: Collision Detection
Reimplemented with official codebase without CD.
Table 7: Detailed results on GraspNet Dataset. Showing APs on RealSense/Kinect split under different friction coefficient.
Backbone Ablation: To further prove the efficiency and robustness of our 2D-convolution-based network, in Table 8, we replace the backbone with the widely used [6, 19, 28] high-dimensional sparse convolution backbone, MinkowskiFCNN [29]. As is shown in the results, our simple 2D convolution network is more efficient than the more complicated sparse convolution network in experiments, especially with high-resolution input. When facing low-resolution input, SparseCNN performs slightly better due to its particular design for sparse input. Significantly, both RNGNet and SparseCNN show quite consistent performance with input patches of different resolutions, which proves the robustness of our region-aware framework.
Backbone Size Seen Similar Novel Time/ms
SparseCNN [29] $64\times 64$ 72.17 61.90 29.85 74
$32\times 32$ 71.76 61.63 29.89 29
$16\times 16$ 68.16 56.68 25.76 18
RNGNet $64\times 64$ 75.20 66.62 32.38 8
$32\times 32$ 72.13 62.81 29.86 5
$16\times 16$ 67.04 55.35 25.63 4
Table 8: Backbone ablation. Showing APs on Realsense split and network inference time.
Grasp Inference with Multi-scale Grippers: As clarified in Section. 2, the Normalized Grasp Space (NGS) demonstrates scale invariance. In other words, our grasp detection model can be easily scaled to accommodate a different range from the fixed gripper width defined in the training dataset without any extra training. In Fig. 9, we illustrate this by enlarging the gripper scale from $w_{gripper}=0.05\ m$ to $w_{gripper}=0.2\ m$, showcasing grasps in a cluttered scene. These results affirm that our framework can seamlessly adapt to grippers of varying sizes. As is shown in Fig. 10, we extend the application of our framework to a much larger indoor scene than the typical table-top settings. The proposed RNGNet successfully generates large-scale grasps for large objects, validating the effectiveness of the Normalized Grasp Space.
To further investigate the effectiveness of our Normalized Grasping Space for object grasping with grippers of different scales, We build scenes with objects with significantly different scales in the Sapien [30] simulator and conduct grasp detection and robot grasping. Experiments prove that our Normalized Grasping Space can generalize well to grippers with different scales and has the potential to apply to industrial scenes, in which data and labels are hard to obtain. Detailed results and visualization can be found in our Supplementary Video.
Figure 9: Visualizations of grasp predictions of different gripper scales. Different colors mean different gripper widths. From (a) to (e), as the gripper scale enlarges, it shows adaptive grasp predictions under different configurations.
Figure 10: Grasp detection in a room-scale scene. (a) Target RGB image and depth image of an ordinary indoor scene with a sofa placed. (b) Extracted patch RGB and normalized z image. (c) Detected grasps of the sofa.
Appendix B Implemetation Details
Normalized Patch Dataset Preparation: Currently, most grasp datasets focus on providing object-level grasp labels [13, 14] or scene-level grasp label [3, 31], which are not directly suitable for our normalized scheme training. Although GraspNet-1Billion offers a large number of RGBD images and grasps annotations, based on the proposed Normalized Grasp Space, we need to construct a patch-based grasp dataset providing normalized patches and grasps annotations. As is shown in Fig. 11, utilizing the scene-level annotations provided in the GraspNet-1Billion dataset, we generate a new patch-based grasp dataset for Regional Normalized Grasp Network training. Firstly we obtain the object segmentation mask in the original dataset and adopt a Gaussian kernel to dilate the mask image for covering possible graspable regions as much as possible. Then a fixed number of patch center candidates are randomly sampled on the dilated mask image and cropped. Also, as is mentioned in [32], depth information is important to generate collision-free grasp detection results. Thus, we add domain randomization to the patch center depth on the RGBD image planes, which provides a patch-based dataset covering more of the possible circumstances and improves the robustness of our training method.
Figure 11: Pipeline of Normalized Patch Dataset Preparation.
Grasping Order and Execution in Clutters: In our cluttered scene clearance experiments, we mainly adopt the perception-planning-grasping open-loop scheme. For each grasp attempt, firstly, the robot moves to the initial home pose and captures the single-view RGBD image, from which multiple normalized grasp regions are extracted. Then, grasps are detected in each region to compose the scene-level grasps for later grasp execution. During the grasping part, the grasp with the highest score is selected as the current target pose. We utilize the MoveIt [33] framework to plan the robot’s trajectory and execute the grasp poses using the UR5e equipped with the Robotiq gripper. RRT [34] serves as our path planner to plan a feasible trajectory to the target pose and execute the final grasp action. To avoid potential collisions and unsafe planning trajectories, MoveIt’s collision box construction tool is employed to define collision geometry. In case of failed grasp execution resulting from collisions or inaccurate grasps, the scene will be restored, and the objects will be re-placed.
Closed-loop Grasping Pipeline: As demostrated in the Algorithm. 1, we integrate grasp detection with simple closed-loop robot controlling and keep tracking the real-time grasp detection results to achieve closed-loop grasping in dynamic scenes.
@@ Appendix B Implemetation Details
1:  ## Tracking phrase
2:  while $\text{dist}(grasps[0].\text{pose},robot.\text{pose})>min\_dis$ do
@@
4:     $centers=\text{uniform\_sample}(M_{workspace})$ sample patch centers within the mask
5:     $patches=\text{normalize}(\text{patch\_extract}(centers,image))$ extract and normalize patch regions
6:     $grasps=\text{RNGNet}(patches)$ detect grasps for objects in the workspace
@@
8:     $robot.\text{set\_speed}(grasps[0].\text{pose}-robot.\text{pose})$ robot moving towards the best grasp pose
9:  end while
10:  ## Grasping phrase
11:  $robot.\text{move\_pose}(grasps[0].\text{pose})$ $robot.\text{grasp}()$ move robot to the final grasp pose and grasp
Hyperparameters and Loss: In the patch extraction, we extract normalized patches with size $S\times S=64\times 64$ by default. We process the input image as the same size $640\times 360$ and adopt the pre-trained Grasp Heatmap Model checkpoint in [15] for a fair performance comparison. In the feature extraction part, we directly adopt the ResNet-18 [35] with half width as the mainstream of our backbone in RNGNet. For the implementation of the Point-wise Gate module, a 3-layer shared MLP with the same width as the corresponding feature maps is adopted to generate gate value for feature maps of each resolution. In our rotation heatmap predictor, the settings about rotation anchor are kept unchanged from those in the [15]. AdamW optimizer is adopted to train our RNGNet.
$L=L_{\theta\_cls}+L_{\theta\_reg}+L_{\gamma\beta}+L_{t}+L_{w},$ (7)
Appendix C Limitations